If a,b,c are real numbers and a+b+c=0, What's a(b-c)^3 + b(c-a)^3 + c(a-b)^3 ?
1
Expert's answer
2019-12-31T06:54:36-0500

"a(b-c)^3+b(c-a)^3+c(a-b)^3="

"= a(b^3\u22123b^2c+3bc^2\u2212c^3)+b(c^3\u22123c^2a+3ca^2\u2212a^3)+c(a^3\u22123a^2b+3ab^2\u2212b^3)="

"=ab ^3 \u22123ab ^2 c+3abc ^2 \u2212ac ^3 + bc ^3 \u22123abc ^2 +3a ^2 bc\u2212ba ^3 +ca ^3 \u22123a ^2 bc+3ab ^2 c\u2212cb ^3\n =a b ^3 \u2212ac ^3 +bc ^3 \u2212ba ^3 +ca ^3 \u2212cb ^3"


"a+b+c=0\u21d2c=\u2212a\u2212b"


"ab ^3 \u2212ac ^3 +bc ^3 \u2212ba ^3 +ca ^3 \u2212cb ^3=ab ^3 \u2212ba ^3 +(\u2212a+b)c ^3 +c(a ^3 \u2212b ^3 )="

"=a b ^3 \u2212ba ^3 +(\u2212a+b)(\u2212a\u2212b) ^3 +(\u2212a\u2212b)(a ^3 \u2212b ^3 )="

"= ab ^3 \u2212ba ^3 +(\u2212a+b)(\u2212a ^3 \u22123a ^2 b\u22123ab ^2 \u2212b ^3 )\u2212a ^4 +ab ^3 \u2212ba ^3 +b ^4="

"=-a ^4 +2ab ^3 \u22122ba ^3 +b ^4 +(a ^4 +3a ^3 b+3a ^2 b ^2 +ab ^3 \u2212ba ^3 \u22123a ^2 b^2 \u22123ab ^3 \u2212b ^4 )=0"

"a(b\u2212c) ^3\n +b(c\u2212a) ^3\n +c(a\u2212b) ^3\n =0"